OpenStudy (anonymous):

simplify 5 sqrt 112 9 sqrt 7 4 sqrt 7 80 sqrt 7 20 sqrt 7

OpenStudy (angela210793):

=5sqr(16*7)=5*4sqr7=20sqr7